A Bibliography of Publications of Jorge Moré

Jorge Moré

Mathematics and Computer Science Division Argonne National Laboratory, 9700 S. Cass Ave.,

Argonne, IL 60439-4844 USA

Tel: +1 708 252 7238 FAX: +1 708 252 5986

E-mail: more@mcs.anl.gov (Internet)

28 November 2023 Version 1.18

Abstract

This bibliography records publications of Jorge Moré.

Title word cross-reference

ϵ [MW96]. n [MR73]. P [MR73]. S [MR73].

-dimensional [MR73]. -functions [MR73]. -optimal [MW96].

1 [GHM80, MGH80]. 1977 [Wat78]. 1981 [Wat82]. 1982 [BGK83]. 1989 [CL90]. 1996 [BI97].

2 [ACM91, AM91, ACMX92].

566 [AFS94].

87 [NF^{+}87].

9th [Wat82].

Active [BM88]. Algorithm [CGM84a, CGM85a, MC80, MGH81a, Mor78, AFS94]. Algorithms [MT89, Mor83, Mor90b, MT92, MT94, MW09]. Analysis [Gol84, Wat78, Wat82]. application [DM74, Mor72]. Applications [NF^{+}87]. Approximation [BI97]. architectures [AM93, AM94]. Art [BGK83]. August [BGK83]. Australia [NF^{+}87]. automatic [AMB^{+}94, BBKM97].

Benchmarking [DM02, MW09]. Biennial [Wat78, Wat82]. Bonn [BGK83]. bound [LM99b, Mor88a, MT89, Mor90b, MT91]. bound-constrained [LM99b]. Bounds [CM87b]. BRENTM [MC80]. Broyden [MT76].

C5 [MGH81a]. Cambridge [BI97]. characterization [DM74]. Cholesky [LM99a]. Class [Mor70]. Classes [Mor74a, MR73]. CO [AG90]. Cody [MM09]. Coercivity [Mor74b]. collection [ACM91, AM91, ACMX92, Mor90a]. Collins [AG90]. coloring [CM83, CM84]. Complementarity [Mor74b, GM97a, Mor74a, Mor96]. Computational [AG90, NF^{+}87, MW11]. Computing [AMB^{+}94, BBKM97, MS83, MM04]. concave [MV90]. Conditions [Mor74b, Mor74a]. Conference [Fos79, NF^{+}87, Wat78, Wat82, BI97]. constrained [CM87a, LM99b, Mor88a, MT89, Mor90b]. Constraints [BM88, BMT90, BM92, BM94, MT91]. continuation [MW96, MW97a]. Convergence [BMT90, Mor70, BDM73, DM74, Mor71, MT76]. convex [BMT90]. Cornell [CL90]. CTAC [NF^{+}87]. CTAC-87 [NF^{+}87]. curvature [MS79].

D [BI97]. decrease [MT92, MT94]. derivative [MW09]. derivative-free [MW09]. derivatives [MW12]. design [Mor80]. Development [Cow84]. developments [Mor83]. diagonal [Mor72]. differentiation [AMB^{+}94, BBKM97]. dimensional [MR73]. directions [MS79]. Distance [MW99, MW96, MW97a]. dominance [Mor72]. Dundee [Wat78, Wat82].

E4 [MGH81a]. environments [GM97b, GM97c]. Equations [AG90, MC79, MC80]. Estimating [CGM84a, CGM84b, CGM85b, MW11, MW12, CGM85a]. Estimation [CM83, CM84]. Evaluation [AM93, AM94, Fos79]. Exposing [BM92, BM94].

F5 [MC80]. factorizations [LM99a]. feasibility [Mor74a]. Fort [AG90].

FORTRAN

[CGM85a, MGH81a, CGM84a, MC80]. free [MW09]. Functions [Mor70, MR73, Mor74a].

Gauss [Mor70, Mor71, Mor70, Mor72]. Generalizations [Mor93, Mor72]. geometry [MW96, MW97a, MW99]. Global [Mor71, Mor96, MW97a, MT76, MW96, MW97b]. Gradient [Mor89, CM87a]. gradients [BBKM97, Mor88b]. graph [CM83, CM84]. guaranteed [MT92, MT94]. Guide [AM91, GHM80, MGH80, MW93].

Held [Wat78, BGK83, NF^{+}87, Wat82]. Hessian [CM84, CGM85a, CGM85b].

Identification [BM88]. IFIP [Fos79]. Impact [BM97]. Implementation [GHM80, Mor79, Mor78]. Incomplete [LM99a]. International [NF^{+}87, BGK83]. Issues [MW97b]. Iterations [Mor70, Mor72]. Ithaca [CL90].

J [BI97]. Jacobian

[AMB^{+}94, CM83, CGM84a, CGM84b]. Jim [MM09]. July [BI97, Wat78]. June [Wat78, Wat82].

knapsack [MV90].

Large [CL90, AM93, AMB^{+}94, AM94, BBKM97, BM97, LM99b, Mor89, Mor90b, MT91, MW97b]. Large-Scale [CL90, AM93, AM94, BBKM97, BM97, Mor89, Mor90b, MW97b]. Levenberg [Mor78]. libraries [MM09]. limited [LM99a]. Line [MT92, MT94]. linear [BMT90]. linearly [CM87a]. local [BDM73].

M. [BI97]. mappings [MR73]. Marquardt [Mor78]. Mathematical [BGK83, Cow84, MM09]. Mathematics [BI97]. Matrices [CGM84a, CGM84b, CGM85b, AMB^{+}94, CM83, CM84, CGM85a]. matrix [Mor72]. measures [DMM06]. memory [LM99a]. Method [MS84, LM99b, MT76, MS79]. Methods [DM77, BDM73, BMT90, CM87a, DM74, DM83, Mor71, Mor83, Mor96]. MINPACK [ACM91, AM91, ACMX92, GHM80, MGH80, MSHG84]. MINPACK-1 [GHM80, MGH80]. MINPACK-2 [ACM91, AM91, ACMX92]. mixed [GM97a]. model [Mor90a]. modified [MS79]. molecular [MW97b]. Motivation [DM77, DM83]. mountain [MM04].

negative [MS79]. NEOS [GM97b, GM97c]. Newton [BDM73, CM87b, DM74, DM77, DM83, LM99b, Mor70, Mor71, MS79, MS84]. noise [MW11]. noisy [MW12]. Nonlinear [AG90, Mor70, Mor72, Mor74b, MC79, MC80, Pow82, MR73, Mor74a, Mor90a, Mor96]. Notes [Mor82]. Numerical [BI97, CL90, Gol84, MC79, MC80, Mor88a, Fos79, Wat78, Wat82].

October [CL90]. optimal [MW96]. Optimality [DMM06]. Optimization [BI97, CL90, GM97b, GM97c, MGH81a, MGH81b, MW93, Pow82, AM93, AM94, BBKM97, BM97, DM02, LM99b, Mor79, Mor80, Mor82, Mor89, MW97b, MW99, MW09].

parallel [AM93, AM94]. partial [BM97]. passes [MM04]. Performance [Fos79, DM02, DMM06, Mor90b, Fos79]. Pioneer [MM09]. Powell [BI97]. Preliminary [ACM91]. problem [ACM91, AM91, ACMX92, Mor93]. Problems [Mor74b, AM93, AM94, CM87a, CM83, CM84, GM97a, LM99b, Mor74a, Mor88a, MT89, Mor89, Mor90a, Mor90b, MV90, MT91, Mor96, MW96, MW97a]. Proceedings [NF^{+}87, Wat78, Wat82, BGK83, CL90, Fos79]. profiles [DM02, DMM06]. Programming [BGK83, MT89, MT91]. Project [MSHG84]. Projected [CM87a, Mor88b]. projection [Mor89]. properties [BMT90]. protein [MW99].

quadratic [MT89, MT91]. Quasi

[CM87b, DM77, DM83, BDM73, DM74].

Quasi-Newton

[CM87b, DM77, DM83, BDM73, DM74].

region [BMT90, MS83, Mor83, Mor93]. regions [Mor88b]. related [MR73]. Remark [AFS94].

Scale [CL90, AM93, AM94, BBKM97, BM97, Mor89, Mor90b, MW97b]. Scotland [Wat82]. search [MT92, MT94]. Seidel [Mor70, Mor71, Mor72]. separability [BM97]. Server [GM97b, GM97c]. simulations [MW12]. Smoothing [GM97a]. Software [CGM84b, CGM85b, Cow84, MGH81a, MGH81b, MW93, DM02, Fos79, Mor79, Mor80, Mor82, Mor83, MM09, Fos79]. Solution [AG90, MC79, MC80, Mor88a, MV90, MT91]. solutions [MW96]. Sources [Cow84]. Sparse [CGM84a, CGM84b, CGM85b, AMB^{+}94, CM83, CM84, CGM85a]. State [BGK83]. states [MM04]. step [MS83]. structures [MW99]. Studies [Gol84]. Subroutine [MC80]. Subroutines [CGM84a, MGH81a, CGM85a]. sufficient [MT92, MT94]. superlinear [BDM73, DM74]. Sydney [NF^{+}87]. symposium [BGK83]. Systems [AG90].

TC [Fos79]. Techniques [NF^{+}87, Mor89]. test [ACM91, AM91, ACMX92]. Testing [MGH81a, MGH81b, Mor79]. Theory [BI97, DM77, DM83, Mor78]. transition [MM04]. tributes [BI97]. Trust [Mor88b, BMT90, MS83, Mor83, Mor93].

Unconstrained [MGH81a, MGH81b]. University [BGK83, CL90, NF^{+}87]. Updates [CM87b]. use [MS79]. User [AM91, MGH80]. using [AMB^{+}94, BBKM97].

vector [AM93, AM94]. Version [ACM91]. via [MW96].

Working [Fos79]. Workshop [CL90].

XI [BGK83].

York [CL90].

References

Averick:1991:MTP

[ACM91] B. M. Averick, R. G. Carter, and J. J. Moré. The MINPACK-2 test problem collection (Preliminary Version). Technical Memorandum ANL/MCS-TM-150, Argonne National Laboratory, Argonne, IL, USA, 1991. Also issued as Preprint 91-76 of the Army High Performance Computing Research Center at the University of Minnesota.

Averick:1992:MTP

[ACMX92] B. M. Averick, R. G. Carter, J. J. Moré, and G-L. Xue. The MINPACK-2 test problem collection. Preprint MCS-P153-0692, Argonne National Laboratory, Argonne, IL, USA, 1992.

Averbukh:1994:RA

[AFS94] Victoria Z. Averbukh, Samuel Figueroa, and Tamar Schlick. Remark on Algorithm 566. ACM Transactions on Mathematical Software, 20(3):282–285, September 1994. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http:/

/www.acm.org/pubs/citations/
journals/toms/1994-20-3/p282-

averbukh/. See [MGH81a].

Allgower:1990:CSN

[AG90] E. L. Allgower and K. Georg, editors. Computational Solution of Nonlinear Systems of Equations (Fort Collins, CO, 1988), volume 26 of Lectures in Applied Mathematics. American Mathematical Society, Providence, RI, USA, 1990. ISBN 0-8218-1131-2. LCCN QA372 .C6374 1990.

Averick:1991:UGM

[AM91] B. M. Averick and J. J. Moré. User guide for the MINPACK-2 test problem collection. Technical Memorandum ANL/MCS-TM-157, Argonne National Laboratory, Argonne, IL, USA, 1991. Also issued as Preprint 91-101 of the Army High Performance Computing Research Center at the University of Minnesota.

Averick:1993:ELS

[AM93] B. M. Averick and J. J. Moré. Evaluation of large-scale optimization problems on vector and parallel architectures. Preprint MCS-P379-0893, Argonne National Laboratory, Argonne, IL, USA, 1993.

Averick:1994:ELS

[AM94] Brett M. Averick and Jorge J. Moré. Evaluation of large-scale optimization problems on vector and parallel architectures. SIAM Journal on Optimization, 4(4): 708–721, 1994. CODEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic).

Averick:1994:CLS

[AMB^{+}94] Brett M. Averick, Jorge J. Moré, Christian H. Bischof, Alan Carle, and Andreas Griewank. Computing large sparse Jacobian matrices using automatic differentiation. SIAM Journal on Scientific Computing, 15(2):285–294, 1994. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Bischof:1997:CGL

[BBKM97] Christian H. Bischof, Ali Bouaricha, Peyvand M. Khademi, and Jorge J. Moré. Computing gradients in large-scale optimization using automatic differentiation. IN-FORMS Journal on Computing, 9 (2):185–194, Spring 1997. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL

https://pubsonline.informs.
org/doi/abs/10.1287/ijoc.9.
2.185.

Broyden:1973:LSC

[BDM73] C. G. Broyden, J. E. Dennis, Jr., and Jorge J. Moré. On the local and superlinear convergence of quasi-Newton methods. Journal of the Institute of Mathematics and its Applications, 12:223–245, 1973. CODEN JMTAA8. ISSN 0020-2932.

Bachem:1983:MPS

[BGK83] Achim Bachem, Martin Grötschel, and B. Korte, editors. Mathematical Programming: the State of the Art [proceedings of the XI international symposium on mathematical programming held at the University of Bonn, from August 23 to 27, Bonn 1982]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. ISBN 0-387-12082-3 (New York), 3-540-12082-3 (Berlin). LCCN QA402.5 .M3529 1983.

Buhmann:1997:ATO

[BI97] Martin D. Buhmann and Arieh Iserles, editors. Approximation Theory and Optimization: tributes to M. J. D. Powell [Conference on Numerical Mathematics in Cambridge, on 27–30 July 1996]. Cambridge University Press, Cambridge, UK, 1997. ISBN 0-521-58190-7 (hardcover). LCCN QA221 .A655 1997.

Burke:1988:IAC

[BM88] James V. Burke and Jorge J. Moré. On the identification of active constraints. SIAM Journal on Numerical Analysis, 25(5): 1197–1211, October 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Burke:1992:EC

[BM92] J. V. Burke and J. J. Moré. Exposing constraints. Preprint MCS-P308-0592, Argonne National Laboratory, Argonne, IL, USA, 1992.

Burke:1994:EC

[BM94] James V. Burke and Jorge J.

Moré. Exposing constraints. SIAM Journal on Optimization, 4(3):573–595, 1994. CODEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic).

Bouaricha:1997:IPS

[BM97] Ali Bouaricha and Jorge J. Moré.

Impact of partial separability on large-scale optimization. Computational Optimization and Applications, 7(1):27–40, 1997. CODEN CPPPEF. ISSN 0926-6003 (print), 1573-2894 (electronic). Computational issues in high performance software for nonlinear optimization (Capri, 1995).

Burke:1990:CPT

[BMT90] James V. Burke, Jorge J. Moré, and Gerardo Toraldo. Convergence properties of trust region methods for linear and convex constraints. Mathematical Programming, 47(3):305–336, 1990. CO-DEN MHPGA4. ISSN 0025-5610 (print), 1436-4646 (electronic).

Coleman:1984:AFS

[CGM84a] Thomas F. Coleman, Burton S.

Garbow, and Jorge J. Moré. Algorithm 618: Fortran subroutines for estimating sparse Jacobian matrices. ACM Transactions on Mathematical Software, 10(3):346–347, September 1984. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Coleman:1984:SES

[CGM84b] Thomas F. Coleman, Burton S. Garbow, and Jorge J. Moré. Software for estimating sparse Jacobian matrices. ACM Transactions on Mathematical Software, 10(3): 329–345, September 1984. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Coleman:1985:AFS

[CGM85a] Thomas F. Coleman, Burton S. Garbow, and Jorge J. Moré. Algorithm 636: FORTRAN subroutines for estimating sparse Hessian matrices. ACM Transactions on Mathematical Software, 11(4):378, December 1985. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

tronic). URL http://www.acm.
org/pubs/citations/journals/
toms/1985-11-4/p378-coleman/

. The title of this paper incorrectly said Algorithm 649; it should be Algorithm 636.

Coleman:1985:SES

[CGM85b] Thomas F. Coleman, Burton S. Garbow, and Jorge J. Moré. Software for estimating sparse Hessian matrices. ACM Transactions on Mathematical Software, 11(4):363–377, December 1985. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

tronic). URL http://www.acm.
org/pubs/citations/journals/
toms/1985-11-4/p363-coleman/

.

Coleman:1990:LSN

[CL90] Thomas Frederick Coleman and Yuying Li, editors. Large-Scale Numerical Optimization: [proceedings of the Workshop on Large Scale Numerical Optimization, Cornell University. Ithaca, New York, October 19–20, 1989]. SIAM, Philadelphia, PA, USA, 1990. ISBN 0-89871-268-8 (paperback). LCCN QA402.5 .L35 1990.

Coleman:1983:ESJ

[CM83] Thomas F. Coleman and Jorge J.

Moré. Estimation of sparse Jacobian matrices and graph coloring problems. SIAM Journal on Numerical Analysis, 20(1):187– 209, February 1983. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Coleman:1984:ESH

[CM84] Thomas F. Coleman and Jorge J.

Moré. Estimation of sparse Hessian matrices and graph coloring problems. Mathematical Programming, 28(3):243–270, 1984. CO-DEN MHPGA4. ISSN 0025-5610 (print), 1436-4646 (electronic).

Calamai:1987:PGM

[CM87a] Paul H. Calamai and Jorge J.

Moré. Projected gradient methods for linearly constrained problems. Mathematical Programming, 39(1):93–116, 1987. CODEN MH-PGA4. ISSN 0025-5610 (print), 1436-4646 (electronic).

Calamai:1987:QNU

[CM87b] Paul H. Calamai and Jorge J.

Moré. Quasi-Newton updates with bounds. SIAM Journal on Numerical Analysis, 24(6):1434– 1441, December 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cowell:1984:SDM

[Cow84] Wayne R. Cowell, editor. Sources and Development of Mathematical Software. Prentice-Hall Series in Computational Mathematics, Cleve Moler, Advisor. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1984. ISBN 0-13-823501-5. xii + 404 pp. LCCN QA76.95 .S68 1984.

Dennis:1974:CSC

[DM74] J. E. Dennis, Jr. and Jorge J. Moré. A characterization of superlinear convergence and its application to quasi-Newton methods. Mathematics of Computation, 28: 549–560, 1974. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dennis:1977:QNM

[DM77] J. E. Dennis, Jr. and Jorge J. Moré. Quasi-Newton methods, motivation and theory. SIAM Review, 19(1):46–89, January 1977. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dennis:1983:QNM

[DM83] J. E. Dennis, Jr. and Jorge J. Moré. Quasi-Newton methods, motivation and theory. Applied Mathematics & Mathematics of Computation. Yingyong Shuxue yu Jisuan Shuxue, (1):1–30, 1983.

Translated from the English by Wen Yu Sun.

Dolan:2002:BOS

[DM02] Elizabeth D. Dolan and Jorge J.

Moré. Benchmarking optimization software with performance profiles. Mathematical Programming, 91(2):201–213, 2002. CO-DEN MHPGA4. ISSN 0025-5610 (print), 1436-4646 (electronic).

Dolan:2006:OMP

[DMM06] Elizabeth D. Dolan, Jorge J. Moré, and Todd S. Munson. Optimality measures for performance profiles. SIAM Journal on Optimization, 16(3):891–909, 2006. CO-DEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic).

Fosdick:1979:PEN

[Fos79] Lloyd Dudley Fosdick, editor. Performance evaluation of numerical software: proceedings of the IFIP TC 2.5 Working Conference on Performance Evaluation of Numerical Software. North-Hol-land, Amsterdam, The Netherlands, 1979. ISBN 0-444-85330-8. LCCN QA297 .I18 1978.

Garbow:1980:IGM

[GHM80] B. S. Garbow, K. E. Hillstrom, and J. J. Moré. Implementation guide for MINPACK-1. Report ANL-80-68, Argonne National Laboratory, Argonne, IL, USA, 1980.

Gabriel:1997:SMC

[GM97a] Steven A. Gabriel and Jorge J.

Moré. Smoothing of mixed complementarity problems. In Michael C. Ferris and Jong-Shi Pang, editors, Complementarity and Variational Problems [Proceedings of the International Conference on Complementarity Problems, Baltimore, Maryland, November 1–4, 1995], pages 105– 116. SIAM, Philadelphia, PA, USA, 1997. ISBN 0-89871-391-9. LCCN QA402.5 .C553 1997.

Gropp:1997:OENa

[GM97b] William Gropp and Jorge J. Moré. Optimization environments and the NEOS Server. Preprint CRPC-TR97708, MCS-P654-0397, Argonne National Laboratory, Argonne, IL, USA, March 1997. ii + 16 pp. URL

https://ftp.mcs.anl.gov/pub/
tech_reports/reports/P654.pdf;
 https://www.osti.gov/biblio/
563264;   https://www.osti.
gov/servlets/purl/563264.

Gropp:1997:OENb

[GM97c] William Gropp and Jorge J. Moré. Optimization environments and the NEOS Server. In Buhmann and Iserles [BI97], pages 167–182. ISBN 0-521-58190-7 (hardcover). LCCN QA221 .A655 1997.

Golub:1984:SNA

[Gol84] Gene H. Golub, editor. Studies in Numerical Analysis, volume 24 of Studies in mathematics. Mathematical Association of America, Washington, DC, 1984. ISBN 0-88385-126-1 (v. 1), 0-88385-100-8 (set). x + 415 pp. LCCN QA297 .S83 1984.

Lin:1999:ICF

[LM99a] Chih-Jen Lin and Jorge J. Moré.

Incomplete Cholesky factorizations with limited memory. SIAM Journal on Scientific Computing, 21(1):24–45, 1999. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Lin:1999:NML

[LM99b] Chih-Jen Lin and Jorge J. Moré.

Newton’s method for large boundconstrained optimization problems. SIAM Journal on Optimization, 9(4):1100–1127, 1999. CO-DEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic). Dedicated to John E. Dennis, Jr., on his 60th birthday.

More:1979:NSN

[MC79] Jorge J. Moré and Michel Y. Cosnard. Numerical solution of nonlinear equations. ACM Transactions on Mathematical Software, 5(1):64–85, March 1979. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

More:1980:ABF

[MC80] J. J. Moré and M. Y. Cosnard. Algorithm 554: BRENTM, A Fortran subroutine for the numerical solution of nonlinear equations [F5]. ACM Transactions on Mathematical Software, 6(2):240–251, June 1980. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

More:1980:UGM

[MGH80] J. J. Moré, B. S. Garbow, and K. E. Hillstrom. User guide for MINPACK-1. Report ANL-80-74, Argonne National Laboratory, Argonne, IL, USA, 1980.

More:1981:AFS

[MGH81a] J. J. Moré, B. S. Garbow, and K. E. Hillstrom. Algorithm 566: FORTRAN subroutines for testing unconstrained optimization software [C5 [E4]]. ACM Transactions on Mathematical Software, 7(1):136–140, March 1981. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See also [AFS94].

More:1981:TUO

[MGH81b] Jorge J. Moré, Burton S. Garbow, and Kenneth E. Hillstrom. Testing unconstrained optimization software. ACM Transactions on Mathematical Software, 7(1):17–41, March 1981. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

More:2004:CMP

[MM04] Jorge J. Moré and Todd S. Munson. Computing mountain passes and transition states. Mathematical Programming, 100(1):151–182, 2004. CODEN MHPGA4. ISSN 0025-5610 (print), 1436-4646 (electronic).

More:2009:JCP

[MM09] Jorge Moré and Cleve Moler. Jim Cody: Pioneer in mathematical software libraries. NA Digest postings, June 28, 2009. URL http:

//www.netlib.org/na-digest-
html/09/v09n26.html#1.

More:1970:CNF

[Mor70] Jorge Jesus Moré. A Class of Nonlinear Functions and the Convergence of Gauss–Seidel and Newton–Gauss–Seidel Iterations. Thesis (Ph.D.), University of Maryland, College Park, MD, USA, 1970. 99

pp. URL https://proquest.
com/docview/302552483?.

More:1971:GCN

[Mor71] Jorge J. Moré. Global convergence of Newton–Gauss–Seidel methods. SIAM Journal on Numerical Analysis, 8:325–336, 1971. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

More:1972:NGM

[Mor72] Jorge J. Moré. Nonlinear generalizations of matrix diagonal dominance with application to Gauss– Seidel iterations. SIAM Journal on Numerical Analysis, 9:357–378, 1972. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

More:1974:CFF

[Mor74a] Jorge J. Moré. Classes of functions and feasibility conditions in nonlinear complementarity problems. Mathematical Programming, 6:327–338, 1974. CODEN MH-PGA4. ISSN 0025-5610 (print), 1436-4646 (electronic).

More:1974:CCN

[Mor74b] Jorge J. Moré. Coercivity conditions in nonlinear complementarity problems. SIAM Review, 16 (1):1–16, ???? 1974. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

More:1978:LMA

[Mor78] Jorge J. Moré. The Levenberg– Marquardt algorithm: Implementation and theory. In Watson [Wat78], pages 105–116. CODEN LNMAA2. ISBN 3-540-08538-6 (print), 3-540-35972-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 630; QA1 .L471 v.630.

                                        URL http://link.springer.
                                        com/book/10.1007/BFb0067690;
                                         http://www.springerlink.com/
                                        content/978-3-540-35972-2.

More:1979:ITO

[Mor79] J. J. Moré. Implementation and testing of optimization software. In Fosdick [Fos79], pages 253– 266. ISBN 0-444-85330-8. LCCN QA297 .I18 1978.

More:1980:DOS

[Mor80] J. J. Moré. On the design of optimization software. Quad. Unione Mat. Italiana, 17:85–102, 1980.

More:1982:NOS

[Mor82] Jorge J. Moré. Notes on optimization software. In Powell [Pow82], pages 339–352. ISBN 0-12-563860-4. LCCN QA402.5 .N66 1981.

More:1983:RDA

[Mor83] J. J. Moré. Recent developments in algorithms and software for trust region methods. In Bachem et al. [BGK83], pages 258–287. ISBN 0-387-12082-3 (New York), 3-540-12082-3 (Berlin). LCCN QA402.5 .M3529 1983.

More:1988:NSB

[Mor88a] Jorge J. Moré. Numerical solution of bound constrained problems. In Noye et al. [NF^{+}87], pages 29– 37. ISBN 0-444-70400-0. LCCN QA299.6 .I561 1987.

More:1988:TRP

[Mor88b] Jorge J. Moré. Trust regions and projected gradients. In Masao Iri and Keiji Yajima, editors, Systems Modelling and Optimization: Proceedings of the 13th IFIP Conference, Tokyo, Japan, August 31– September 4, 1987, volume 113 of Lecture Notes in Control and Information Sciences, pages 1– 13. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1988. ISBN 0-387-19238-7, 3-540-19238-7. ISSN 0170-8643 (print), 1610-7411 (electronic). LCCN QA402 .S95823 1988.

More:1989:GPT

[Mor89] Jorge J. Moré. Gradient projection techniques for large-scale optimization problems. In Proceedings of the 28th IEEE Conference on Decision and Control, Vol. 1–3 (Tampa, FL, 1989), pages 378–381. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1989.

More:1990:CNM

[Mor90a] Jorge J. Moré. A collection of nonlinear model problems. In Allgower and Georg [AG90], pages 723–

762. ISBN 0-8218-1131-2. LCCN

QA372 .C6374 1990.

More:1990:PAL

[Mor90b] Jorge J. Moré. On the performance of algorithms for large-scale bound constrained problems. In Coleman and Li [CL90], pages 32–

45. ISBN 0-89871-268-8 (paperback). LCCN QA402.5 .L35 1990.

More:1993:GTR

[Mor93] J. J. Moré. Generalizations of the trust region problem. Optimization Methods and Software, 2:189– 209, 1993. ISSN 1055-6788.

More:1996:GMN

[Mor96] Jorge J. Moré. Global methods for nonlinear complementarity problems. Mathematics of Operations Research, 21(3):589– 614, 1996. CODEN MOREDQ. ISSN 0364-765X (print), 1526-5471 (electronic).

More:1973:FRC

[MR73] J. Moré and W. Rheinboldt. On P - and S-functions and related classes of n-dimensional nonlinear mappings. Linear Algebra and its Applications, 6(??):45–68, ???? 1973. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/0024379573900062.

More:1979:UDN

[MS79] Jorge J. Moré and Danny C. Sorensen. On the use of directions of negative curvature in a modified Newton method. Mathematical Programming, 16(1):1–20, 1979. CODEN MHPGA4. ISSN 0025-5610 (print), 1436-4646 (electronic).

More:1983:CTR

[MS83] Jorge J. Moré and D. C. Sorensen.

Computing a trust region step. SIAM Journal on Scientific and Statistical Computing, 4(3):553– 572, 1983. CODEN SIJCD4. ISSN 0196-5204.

More:1984:NM

[MS84] Jorge J. Moré and D. C. Sorensen.

Newton’s method. In Golub [Gol84], pages 29–82. ISBN 0-88385-126-1 (v. 1), 0-88385-100-8 (set). LCCN QA297 .S83 1984.

More:1984:MP

[MSHG84] J. J. Moré, D. C. Sorensen, K. E.

Hillstrom, and B. S. Garbow. The MINPACK project. In Cowell [Cow84], pages 88–111 (of xii + 404). ISBN 0-13-823501-5. LCCN QA76.95 .S68 1984.

More:1976:GCB

[MT76] J. J. Moré and J. A. Trangenstein. On the global convergence of Broyden’s method. Mathematics of Computation, 30(135):523–540, 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

More:1989:ABC

[MT89] Jorge J. Moré and Gerardo Toraldo. Algorithms for bound constrained quadratic programming problems. Numerische Mathematik, 55(4):377–400, 1989. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

More:1991:SLQ

[MT91] Jorge J. Moré and Gerardo Toraldo. On the solution of large quadratic programming problems with bound constraints. SIAM Journal on Optimization, 1(1):93– 113, 1991. CODEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic).

More:1992:LSA

[MT92] J. J. Moré and D. J. Thuente. Line search algorithms with guaranteed sufficient decrease. Preprint MCS-P330-1092, Argonne National Laboratory, Argonne, IL, USA, 1992.

More:1994:LSA

[MT94] Jorge J. Moré and David J. Thuente. Line search algorithms with guaranteed sufficient decrease. ACM Transactions on Mathematical Software, 20(3): 286–307, 1994. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

More:1990:SCK

[MV90] Jorge J. Moré and Stephen A. Vavasis. On the solution of concave knapsack problems. Mathematical Programming, 49(3): 397–411, November 1990. CO-DEN MHPGA4. ISSN 0025-5610 (print), 1436-4646 (electronic).

More:1993:OSG

[MW93] Jorge J. Moré and Stephen J.

Wright. Optimization Software Guide, volume 14 of Frontiers in Applied Mathematics. SIAM, Philadelphia, PA, USA, 1993. ISBN 0-89871-322-6. xii + 154 pp. LCCN QA402.5 .M67 1993.

More:1996:OSD

[MW96] Jorge J. Moré and Zhijun Wu. ϵoptimal solutions to distance geometry problems via global continuation. In P. M. Pardalos, D. Shalloway, and G. Xue, editors, Global Minimization of Nonconvex Energy Functions: Molecular Conformation and Protein Folding [DI-MACS Workshop, March 20–21, 1995, New Brunswick, NJ], volume 23 of DIMACS series in discrete mathematics and theoretical computer science, pages 151–168. American Mathematical Society, Providence, RI, USA, 1996. ISBN 0-8218-0471-5. ISSN 1052-1798.

More:1997:GCD

[MW97a] Jorge J. Moré and Zhijun Wu.

Global continuation for distance geometry problems. SIAM Journal on Optimization, 7(3):814– 836, 1997. CODEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic).

More:1997:ILS

[MW97b] Jorge J. Moré and Zhijun Wu. Issues in large-scale global molecular optimization. In Lorenz T. Biegler, Thomas F. Coleman, Andrew R. Conn, and Fadil N. Santosa, editors, Large-scale Optimization with Applications, Part III: Molecular Structure and Optimization (Minneapolis, MN, 1995), volume 94 of IMA Volumes in Mathematics and its Applications, pages 99–121. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 0-387-98288-4.

More:1999:DGO

[MW99] Jorge J. Moré and Zhijun Wu. Distance geometry optimization for protein structures. Journal of Global Optimization, 15(3):219– 234, 1999. CODEN JGOPEO. ISSN 0925-5001 (print), 1573-2916 (electronic).

More:2009:BDF

[MW09] Jorge J. Moré and Stefan M. Wild. Benchmarking derivative-free optimization algorithms. SIAM Journal on Optimization, 20(1): 172–191, 2009. CODEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic).

More:2011:ECN

[MW11] Jorge J. Moré and Stefan M. Wild. Estimating computational noise. SIAM Journal on Scientific Computing, 33(3):1292–1314, 2011. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

More:2012:EDN

[MW12] Jorge J. Moré and Stefan M. Wild. Estimating derivatives of noisy simulations. ACM Transactions on Mathematical Software, 38(3):Art. 19, 21, 2012. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Noye:1987:CTA

[NF^{+}87] John Noye, Clive Fletcher, et al., editors. Computational Techniques and Applications, CTAC-87: Proceedings of the 1987 International Conference on Computational Techniques and Applications, held at the University of Sydney, Australia. North-Hol-land, Amsterdam, The Netherlands, 1987. ISBN 0-444-70400-0. LCCN QA299.6 .I561 1987.

Powell:1982:NO

[Pow82] M. J. D. Powell, editor. Nonlinear Optimization 1981, NATO Conference Series II: Systems Science. Academic Press, New York, NY, USA, 1982. ISBN 0-12-563860-4. LCCN QA402.5 .N66 1981.

Watson:1978:NAP

[Wat78] G. A. Watson, editor. Numerical Analysis: Proceedings of the Biennial Conference Held at Dundee, June 28–July 1, 1977, volume 630 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNMAA2. ISBN 3-540-08538-6 (print), 3-540-35972-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 630; QA1 .L471 v.630.

URL http://link.springer.
com/book/10.1007/BFb0067690;
 http://www.springerlink.com/
content/978-3-540-35972-2.

Watson:1982:NAP

[Wat82] George Alistair Watson, editor. Numerical Analysis: Proceedings of the 9th Biennial Conference, held at Dundee, Scotland, June 23–26, 1981, volume 912 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. CO-DEN LNMAA2. ISBN 0-387-11199-9 (softcover), 3-540-11199-9 (softcover), 3-540-39009-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 912; QA1 .L471; QA297 .D915n 1981. URL

                                        http://www.springerlink.com/
                                        content/978-3-540-39009-1.